### v1

@@ 2 The basic JDSVD method
Notice that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}$. Bound (3.14) follows from applying (3.12) to the above min-max problem.
Case (iii): If $(\sigma_{\min}+\sigma_{\min,2})/2<\tau<(\sigma_{\max}+\sigma_{\max,2})/2$, denote by $\sigma_{\rm L}$ and $\sigma_{\rm R}$ respectively the singular values of $A$ closest to $\tau$ from the left and right among $\sigma_{2},\dots,\sigma_{N}$. In this case, the $\lambda_{i}$ lie in
@@
(3.20) $\subset[-(\sigma_{\max}+\tau),-|\tau-\sigma_{2}|]\cup[|\tau-\sigma_{2}|,\sigma_{\max}+\tau].$
Notice that the two intervals in (3.20) have the same lengths $\sigma_{\max}+\tau-|\tau-\sigma_{2}|$. Then (3.14) follows from applying (3.12) to the interval union in (3.20). ∎
Remark 1.
The convergence rates of MINRES, respectively, depend on
(3.21) $\gamma_{1}=\frac{\tau-\sigma_{\max,2}}{\tau+\sigma_{\max,2}},\qquad\gamma_{2}=\sqrt{\frac{\sigma_{\min,2}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}}},\qquad\gamma_{3}=\frac{|\tau-\sigma_{2}|}{\sigma_{\max}+\tau}.$
The better $\sigma_{\max,2}$, $\sigma_{\min,2}$ or $\sigma_{2}$ is separated from $\tau$, the smaller the convergence rate $1-\frac{2\sqrt{\gamma_{1}}}{1+\sqrt{\gamma_{1}}}$, $1-\frac{2\gamma_{2}}{1+\gamma_{2}}$ or $1-\frac{2\gamma_{3}}{1+\gamma_{3}}$ in (3.13)–(3.15) is, so that MINRES converges faster.
Remark 2.
For $\sigma_{1}=\sigma_{\max}$ or $\sigma_{\min}$, which correspond to $\tau\approx\sigma_{\max}$ or $\sigma_{\min}$, we have
$\gamma_{1}\approx\frac{\sigma_{\max}-\sigma_{\max,2}}{\sigma_{\max}+\sigma_{\max,2}}\qquad\mbox{and}\qquad\gamma_{2}\approx\sqrt{\frac{\sigma_{\min,2}^{2}-\sigma_{\min}^{2}}{\sigma_{\max}^{2}-\sigma_{\min}^{2}}}.$
Suppose $\sigma_{\max}\gg\sigma_{\min}$, $\sigma_{\max,2}\approx\sigma_{\max}$ and $\sigma_{\min,2}\approx\sigma_{\min}$. Then
$\frac{\sqrt{\gamma_{1}}}{\gamma_{2}}\approx\sqrt{\frac{\sigma_{\max}(\sigma_{\max}-\sigma_{\max,2})}{2(\sigma_{\min,2}-\sigma_{\min})(\sigma_{\min,2}+\sigma_{\min})}}\approx\sqrt{\frac{\sigma_{\max}}{2\sigma_{\min}}}>1$
if $\sigma_{\max}-\sigma_{\max,2}\approx\sigma_{\min,2}-\sigma_{\min}$, and the ratio is bigger if $\sigma_{\max}-\sigma_{\max,2}>\sigma_{\min,2}-\sigma_{\min}$ considerably, which is often true in applications. For an interior $\sigma_{1}\approx\tau$, we have
$\gamma_{3}\approx\frac{|\sigma_{1}-\sigma_{2}|}{\sigma_{\max}+\sigma_{1}},$
which is generally comparable to $\gamma_{1}$ in size. Since $\gamma_{1},\gamma_{3}<1$, we generally have $\sqrt{\gamma_{1}}>\gamma_{3}$ considerably. In summary, for $\sigma_{1}=\sigma_{\max}$, the linear system (2.7) is negative definite with $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$, and MINRES is generally faster than it is for the indefinite system (2.7) when $\sigma_{1}=\sigma_{\min}$ or an interior singular value.
Starting with Theorem 3.3, we are able to analyze the convergence of MINRES when $\tilde{u}_{1}\approx u_{1}$ and $\tilde{v}_{1}\approx v_{1}$. First, we show that $\widetilde{B}_{1}^{\prime}$ in (3.3) is orthogonally similar to $B_{1}^{\prime}$ in (3.17) with squared errors of $\tilde{u}_{1}$ and $\tilde{v}_{1}$.
@@
(3.22) $\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\|A\|\left(\sin\phi+\sin\psi\right)^{2},$
where $Z$ is some $(M+N-2)$-by-$(M+N-2)$ orthogonal matrix; see (3.31).
Proof.
By (3.16), expand the orthonormal $\widetilde{P}_{1}$ and $\widetilde{Q}_{1}$ in (3.1) as the orthogonal sums
(3.23) $\widetilde{P}_{1}=P_{1}E_{1}+u_{1}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{1}=Q_{1}F_{1}+v_{1}F_{2},$
where $E_{1}\in\mathbb{R}^{(M-1)\times(M-1)}$, $E_{2}\in\mathbb{R}^{1\times(M-1)}$ and $F_{1}\in\mathbb{R}^{(N-1)\times(N-1)}$, $F_{2}\in\mathbb{R}^{1\times(N-1)}$. Then we have
(3.24) $\sin\phi=\|\widetilde{P}_{1}^{T}u_{1}\|=\|E_{2}\|\qquad\mbox{and}\qquad\sin\psi=\|\widetilde{Q}_{1}^{T}v_{1}\|=\|F_{2}\|.$
Let the SVDs of the square matrices $E_{1}$ and $F_{1}$ be
(3.25) $X_{1}^{T}E_{1}X_{2}=\Sigma_{E}\qquad\mbox{and}\qquad Y_{1}^{T}F_{1}Y_{2}=\Sigma_{F}.$
Notice $\|I-\Sigma_{E}\|\leq\|I-\Sigma_{E}^{T}\Sigma_{E}\|$ as $\|\Sigma_{E}\|\leq 1$ and the diagonal $\Sigma_{E}$ is square and nonnegative definite. Making use of (3.25) and $E_{1}^{T}E_{1}+E_{2}^{T}E_{2}=I$, we obtain
(3.26) $\|I-\Sigma_{E}\|\leq\|X_{2}^{T}(I-E_{1}^{T}E_{1})X_{2}\|=\|E_{2}^{T}E_{2}\|=\|E_{2}\|^{2}=\sin^{2}\phi,$
where the last equality holds by (3.24). Likewise,
(3.27) $\|I-\Sigma_{F}\|\leq\sin^{2}\psi.$
Since $P_{1}^{T}Av_{1}=\bm{0}$ and $u_{1}^{T}AQ_{1}=\bm{0}$, it is known from (3.17) and (3.23) that $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=E_{1}^{T}\Sigma_{1}^{\prime}F_{1}+\sigma_{1}E_{2}^{T}F_{2}$. Exploiting (3.25), we have
(3.28) $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}.$
Define
(3.29) $\widehat{\Sigma}_{1}^{\prime}=X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}=X_{2}\Sigma_{1}^{\prime\prime}Y_{2}^{T}\quad\mbox{with}\quad\Sigma_{1}^{\prime\prime}=X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}.$
Subtracting $\widehat{\Sigma}_{1}^{\prime}$ from both hand sides of (3.28) and taking $2$-norms, we obtain
(3.30) $\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\!-\widehat{\Sigma}_{1}^{\prime}\|$ $=$ $\|X_{2}(\Sigma_{E}\Sigma_{1}^{\prime\prime}\Sigma_{F}-\Sigma_{1}^{\prime\prime})Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}\|$
$=$ $\|X_{2}\left((\Sigma_{E}-I)\Sigma_{1}^{\prime\prime}\Sigma_{F}+\Sigma_{1}^{\prime\prime}(\Sigma_{F}-I)\right)Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}\|$
$\leq$ $\|\Sigma_{E}-I\|\|\Sigma_{1}^{\prime\prime}\|+\|\Sigma_{1}^{\prime\prime}\|\|\Sigma_{F}-I\|+\sigma_{1}\|E_{2}\|\|F_{2}\|$
@@
where the second inequality holds by making use of (3.24), (3.26), (3.27) and $\|\Sigma_{1}^{\prime\prime}\|=\|\Sigma_{1}^{\prime}\|\leq\|A\|$, $\sigma_{1}\leq\|A\|$. Define the orthogonal matrix
(3.31) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2)\times(M+N-2)}.$
By the definitions of $\widetilde{B}_{1}^{\prime}$ and $B_{1}^{\prime}$ in (3.3) and (3.17) as well as that of $\widehat{\Sigma}_{1}^{\prime}$ in (3.29), a manipulation shows
$\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|=\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}-X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}\|=\|\widetilde{P}_{1}^{T}\!A\widetilde{Q}_{1}\!-\widehat{\Sigma}_{1}^{\prime}\|.$
Applying (3.30) to the above establishes (3.22). ∎
With the help of Theorems 3.33.4, we can establish the convergence results on MINRES for the correction equation (3.2), i.e., (2.7).
@@
With the notations of Theorems 3.4, suppose that $\tilde{u}_{1}$ and $\tilde{v}_{1}$ satisfy
(3.32) $\delta:=\|A\|(\sin\phi+\sin\psi)^{2}<|\sigma_{2}-\tau|.$
@@
(3.33) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2}+\delta)/(\tau-\sigma_{\max,2}-\delta)}}\right)^{j};$
@@
(3.34) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,2}-\delta)^{2}-\tau^{2}}}}\right)^{\left[\frac{j-j_{o}}{2}\right]};$
@@
(3.35) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{2}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
Proof.
Since the matrix $Z$ in (3.31) is orthogonal, the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ are those of $B_{1}^{\prime}$, i.e., the ones in (3.18). Label the eigenvalues $\lambda_{i}$ and $\tilde{\lambda}_{i}$ of $B_{1}^{\prime}$ and $\widetilde{B}_{1}^{\prime}$ in the ascending or descending order. By the eigenvalue perturbation theory (cf. Theorem 10.3.1 of [21]), we obtain
(3.36) $|\tilde{\lambda}_{i}-\lambda_{i}|\leq\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\delta,$
where the last inequality holds by (3.22) and (3.32).
(i) For $\sigma_{1}=\sigma_{\max}$, since the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ lie in the interval in (3.19), bound (3.36) shows that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
(3.37) $[-(\sigma_{\max,2}+\tau)-\delta,-(\tau-\sigma_{\max,2})+\delta].$
Note that $\sigma_{\max,2}=\sigma_{2}$. Condition (3.32) indicates $\tau-\sigma_{\max,2}>\delta$. Then taking $\alpha=\tau-\sigma_{\max,2}-\delta$ and $\beta=\tau+\sigma_{\max,2}+\delta$ in (3.11) of Theorem 3.2 proves (3.33).
(ii) Since that $\widetilde{B}_{1}^{\prime}$ always has an $(M-N)$-multiple eigenvalue $\tilde{\lambda}=-\tau$ for $M>N$, for $\sigma_{1}=\sigma_{\min}$, (3.18) and (3.36) indicate that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
@@
Notice that $\sigma_{\min,2}=\sigma_{2}$ and (3.32) mean that $\sigma_{\min,2}-\tau>\delta$, and observe that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}+2\delta$. Following the proof of item (ii) of Theorem 3.3, we obtain bound (3.34).
(iii) For $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$, since the eigenvalues of $Z^{T}B_{1}^{\prime}Z$ lie in the interval of (3.20), bound (3.36) means that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
(3.38) $[-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]\cup[|\sigma_{2}-\tau|-\delta,\sigma_{\max}-\tau+\delta]$
$\subset[-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]\cup[|\sigma_{2}-\tau|-\delta,\sigma_{\max}+\tau+\delta]$ ,
Since the two intervals in (3.38) have the equal length $\sigma_{\max}+\tau-|\sigma_{2}-\tau|+2\delta$, applying the proof of item (ii) of Theorem 3.2 to this interval union establishes (3.35). ∎
Remark 3.
It can be verified that the factors in (3.33)–(3.35) are approximations to $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (3.21) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,2}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,2}-\tau})$ or $\mathcal{O}(\frac{\delta}{|\sigma_{2}-\tau|})$, respectively. Therefore, for $\frac{\delta}{\tau-\sigma_{\max,2}}$, $\frac{\delta}{\sigma_{\min,2}-\tau}$ or $\frac{\delta}{|\sigma_{2}-\tau|}$ fairly small, Remarks 12 on Theorem 3.3 apply to Theorem 3.5.

### v2

@@ 3 Convergence analysis on MINRES for solving (9)
Notice that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}$. Bound (23) follows from applying (21) to the above min-max problem.
Case (iii): In this case, denote by $\sigma_{\rm L}$ and $\sigma_{\rm R}$ the singular values of $A$ closest to $\tau$ from the left and right among $\sigma_{2},\dots,\sigma_{N}$, respectively. Then the $\lambda_{i}$ lie in
@@
(29) $\subset[-(\sigma_{\max}+\tau),-|\tau-\sigma_{2}|]\cup[|\tau-\sigma_{2}|,\sigma_{\max}+\tau].$
Notice that the two intervals in (29) have the same lengths $\sigma_{\max}+\tau-|\tau-\sigma_{2}|$. Then (23) follows from applying (21) to the interval union in (29).
Remark 3.6.
The convergence rates of MINRES respectively depend on
(30) $\gamma_{1}=\frac{\tau-\sigma_{\max,2}}{\tau+\sigma_{\max,2}},\qquad\gamma_{2}=\frac{\sigma_{\min,2}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}},\qquad\gamma_{3}=\frac{|\tau-\sigma_{2}|}{\sigma_{\max}+\tau}.$
The better $\sigma_{\max,2}$, $\sigma_{\min,2}$ or $\sigma_{2}$ is separated from $\tau$, the smaller $1-\frac{2\sqrt{\gamma_{1}}}{1+\sqrt{\gamma_{1}}}$, $1-\frac{2\sqrt{\gamma_{2}}}{1+\sqrt{\gamma_{2}}}$ or $1-\frac{2\gamma_{3}}{1+\gamma_{3}}$ in (22)–(24) is, causing that MINRES converges faster.
Remark 3.7.
For $\sigma_{1}=\sigma_{\max}$ or $\sigma_{\min}$, which corresponds to $\tau\approx\sigma_{\max}$ or $\sigma_{\min}$, we have
$\gamma_{1}\approx\frac{\sigma_{\max}-\sigma_{\max,2}}{\sigma_{\max}+\sigma_{\max,2}}\qquad\mbox{or}\qquad\gamma_{2}\approx\frac{\sigma_{\min,2}^{2}-\sigma_{\min}^{2}}{\sigma_{\max}^{2}-\sigma_{\min}^{2}}.$
Suppose $\sigma_{\max,2}\approx\sigma_{\max}$ and $\sigma_{\min,2}\approx\sigma_{\min}$ roughly. Then
$\frac{\gamma_{1}}{\gamma_{2}}\approx\frac{\sigma_{\max}(\sigma_{\max}-\sigma_{\max,2})}{\sigma_{\min}(\sigma_{\min,2}-\sigma_{\min})}\gg 1$
if $\sigma_{\max}-\sigma_{\max,2}\approx\sigma_{\min,2}-\sigma_{\min}$ and $\sigma_{\max}\gg\sigma_{\min}$, and the ratio is even bigger if $\sigma_{\max}-\sigma_{\max,2}>\sigma_{\min,2}-\sigma_{\min}$ considerably, a typical case in applications. As a consequence, MINRES is generally much faster for $\sigma_{1}=\sigma_{\max}$ than it is for $\sigma_{1}=\sigma_{\min}$.
Starting with Theorem 3.4, we are able to analyze the convergence of MINRES when $\tilde{u}_{1}\approx u_{1}$ and $\tilde{v}_{1}\approx v_{1}$. First, we show that $\widetilde{B}_{1}^{\prime}$ in (12) is orthogonally similar to $B_{1}^{\prime}$ in (26) with squared errors of $\tilde{u}_{1}$ and $\tilde{v}_{1}$.
@@
(31) $\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\|A\|\left(\sin\phi+\sin\psi\right)^{2},$
where $Z$ is some $(M+N-2)$-by-$(M+N-2)$ orthogonal matrix; see (39).
Proof 3.9.
By (25), expand the orthonormal $\widetilde{P}_{1}$ and $\widetilde{Q}_{1}$ in (10) as the orthogonal sums
(32) $\widetilde{P}_{1}=P_{1}E_{1}+u_{1}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{1}=Q_{1}F_{1}+v_{1}F_{2},$
where $E_{1}\in\mathbb{R}^{(M-1)\times(M-1)}$, $E_{2}\in\mathbb{R}^{1\times(M-1)}$ and $F_{1}\in\mathbb{R}^{(N-1)\times(N-1)}$, $F_{2}\in\mathbb{R}^{1\times(N-1)}$. Then
(33) $\sin\phi=\|\widetilde{P}_{1}^{T}u_{1}\|=\|E_{2}\|\qquad\mbox{and}\qquad\sin\psi=\|\widetilde{Q}_{1}^{T}v_{1}\|=\|F_{2}\|.$
Let the SVDs of $E_{1}$ and $F_{1}$ be
(34) $X_{1}^{T}E_{1}X_{2}=\Sigma_{E}\qquad\mbox{and}\qquad Y_{1}^{T}F_{1}Y_{2}=\Sigma_{F}.$
Notice $\|I-\Sigma_{E}\|\leq\|I-\Sigma_{E}^{T}\Sigma_{E}\|$ as $\|\Sigma_{E}\|\leq 1$ and the diagonal $\Sigma_{E}$ is square and nonnegative definite. Making use of (34) and $E_{1}^{T}E_{1}+E_{2}^{T}E_{2}=I$, we obtain
(35) $\|I-\Sigma_{E}\|\leq\|X_{2}^{T}(I-E_{1}^{T}E_{1})X_{2}\|=\|E_{2}^{T}E_{2}\|=\|E_{2}\|^{2}=\sin^{2}\phi,$
where the last equality holds by (33). Likewise,
(36) $\|I-\Sigma_{F}\|\leq\sin^{2}\psi.$
Since $P_{1}^{T}Av_{1}=\bm{0}$ and $u_{1}^{T}AQ_{1}=\bm{0}$, it is known from (26) and (32) that $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=E_{1}^{T}\Sigma_{1}^{\prime}F_{1}+\sigma_{1}E_{2}^{T}F_{2}$. Exploiting (34), we have
(37) $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}.$
Note that $X_{i},Y_{i}$, $i=1,2$ are orthogonal matrices. Subtracting $X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}$ from both hand sides of (37) and taking $2$-norms, we obtain
(38) $\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\!-\!X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}\|$ $=$ $\|\Sigma_{E}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}-X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}+\sigma_{1}X_{2}^{T}E_{2}^{T}F_{2}Y_{2}\|$
$\leq$ $\|\!(\Sigma_{E}\!-\!I)X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}\!+\!X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}(\Sigma_{F}\!-\!I)\!\|\!+\!\sigma_{1}\|E_{2}^{T}\!F_{2}\|$
$\leq$ $\|\Sigma_{E}-I\|\|\Sigma_{1}^{\prime}\|+\|\Sigma_{1}^{\prime}\|\|\Sigma_{F}-I\|+\sigma_{1}\|E_{2}\|\|F_{2}\|$
@@
where the third inequality holds by making use of (33), (35), (36) and $\|\Sigma_{1}^{\prime\prime}\|=\|\Sigma_{1}^{\prime}\|\leq\|A\|$, $\sigma_{1}\leq\|A\|$. Define the orthogonal matrix
(39) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2)\times(M+N-2)}.$
By the definitions of $\widetilde{B}_{1}^{\prime}$ and $B_{1}^{\prime}$ in (12) and (26), a manipulation shows
$\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|=\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}-X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}\|.$
Applying (38) to the above establishes (31).
With the help of Theorems 3.4 and 3.8, we can establish the convergence results on MINRES for the correction equation (11), i.e., (9).
@@
With the notations of Theorem 3.8, suppose that $\tilde{u}_{1}$ and $\tilde{v}_{1}$ satisfy
(40) $\delta:=\|A\|(\sin\phi+\sin\psi)^{2}<|\sigma_{2}-\tau|.$
@@
(41) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2}+\delta)/(\tau-\sigma_{\max,2}-\delta)}}\right)^{j};$
@@
(42) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,2}-\delta)^{2}-\tau^{2}}}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]};$
@@
(43) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{2}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
Proof 3.11.
Since the matrix $Z$ in (39) is orthogonal, the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ are those of $B_{1}^{\prime}$, i.e., the ones in (27). Label the eigenvalues $\lambda_{i}$ and $\tilde{\lambda}_{i}$ of $B_{1}^{\prime}$ and $\widetilde{B}_{1}^{\prime}$ in the ascending or descending order. By the eigenvalue perturbation theory (cf. Theorem 10.3.1 of [20]), we obtain
(44) $|\tilde{\lambda}_{i}-\lambda_{i}|\leq\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\delta,$
where the last inequality holds by (31) and (40).
(i) For $\sigma_{1}=\sigma_{\max}$, since the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ lie in the interval in (28), bound (44) shows that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
(45) $[-(\sigma_{\max,2}+\tau)-\delta,-(\tau-\sigma_{\max,2})+\delta].$
Note that $\sigma_{\max,2}=\sigma_{2}$. Condition (40) indicates $\tau-\sigma_{\max,2}>\delta$. Then taking $\alpha=\tau-\sigma_{\max,2}-\delta$ and $\beta=\tau+\sigma_{\max,2}+\delta$ in (20) of Theorem 3.3 proves (41).
(ii) Since $\widetilde{B}_{1}^{\prime}$ has an $(M-N)$-multiple eigenvalue $\tilde{\lambda}=-\tau$ for $M>N$, for $\sigma_{1}=\sigma_{\min}$, (27) and (44) indicate that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
@@
Notice that $\sigma_{\min,2}=\sigma_{2}$ and (40) mean that $\sigma_{\min,2}-\tau>\delta$, and observe that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}+2\delta$. Following the proof of item (ii) of Theorem 3.4, we obtain bound (42).
(iii) For $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$, since the eigenvalues of $Z^{T}B_{1}^{\prime}Z$ lie in the interval defined by (29), bound (44) means that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
$[-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]$$\cup$ $[|\sigma_{2}-\tau|-\delta,\sigma_{\max}-\tau+\delta]$
(46) $\subset\ [-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]$$\cup$ $[|\sigma_{2}-\tau|-\delta,\sigma_{\max}+\tau+\delta],$
Since the two intervals in (46) have the equal length $\sigma_{\max}+\tau-|\sigma_{2}-\tau|+2\delta$, applying the proof of item (iii) of Theorem 3.4 to this interval union establishes (43).
Remark 3.12.
It can be verified that the factors in (41)–(43) are approximations to $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (30) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,2}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,2}-\tau})$ or $\mathcal{O}(\frac{\delta}{|\sigma_{2}-\tau|})$, respectively. Therefore, for $\delta$ reasonably small, Remarks 3.6 and 3.7 apply to Theorem 3.10.